OpenStudy (anonymous):

Find the distance between these points. A (5, 8), B(-3, 4)

OpenStudy (anonymous):

AB=?

OpenStudy (anonymous):

use the formula

OpenStudy (anonymous):

\[\Large AB=\sqrt{(x_A-x_B)^2+(y_A-y_B)^2}\]

OpenStudy (anonymous):

do the formula ( square root of y2-y1 + x2-x1) and then solve. Or graph it and use pythagorm theory

OpenStudy (lopezking1):

\[4\sqrt{5}\] @emoteddybearxx

OpenStudy (anonymous):

d = √(5+3)^2 + (8-4)^2 d = √(8)^2 + (4)^2 d = √64+16 d = √80 d = 4√5

OpenStudy (anonymous):

omg thank you.

OpenStudy (lopezking1):

welcome lol